If x is positive show that. x>log(1+x)>x-(x^2/2)


1
Expert's answer
2022-01-02T17:36:38-0500

To prove

"x>log(1+x)>x-\frac{x^{2}}{2}" when "x>0"

We first suppose that

"f(x)=x-log(1+x)"

Then

"f^\prime(x)=1-\frac{1}{1+x}"

"\implies" "f^\prime(x)=\frac{x}{1+x}"

"\implies" "f^\prime(x)>0" as "x>0"

"\implies" "f(x)>f(0)"

"\implies" "f(x)>0" as "f(0)=0"

"\implies" "x-log(1+x)>0"

"\implies" "x>log(1+x) ----(1)"


Next, Suppose that


"g(x)=log(1+x)-(x-\frac{x^{2}}{2})"

when "x>0"

"\implies" "g^\prime(x)=\frac{1}{1+x}-(1-x)"

"\implies" "g^\prime(x)=\frac{1}{1+x}-1+x"

"\implies" "g^\prime(x)=\frac{1-1-x}{1+x}+x"

"\implies" "g^\prime(x)=x-\frac{x}{1+x}"

"\implies" "g^\prime(x)=\frac{x^{2}}{1+x}"

"\implies" "g^\prime(x)>0" as "x>0"

So "g(x)" is increasing for "x>0"


"\implies" "g(x)>g(0)"

"\implies" "g(x)>0" as "g(0)=0"

"\implies" "log(1+x)-(x-\frac{x^{2}}{2})>0"

when "x>0"

"\implies" "log(1+x)>(x-\frac{x^{2}}{2})---(2)"


Combining (1) and (2), we have


"x>log(1+x)>x-\frac{x^{2}}{2}"

when "x>0"

Hence proved.